# Markov chain Monte Carlo without evaluating the target: an auxiliary variable approach

### v2

@@ 6 Future directions
Another intriguing question is the source of performance gains in minibatch MCMC algorithms. Empirically, these algorithms often achieve ESS/s values 10-100 times greater than MCMC algorithms running on the entire dataset. However, it is proven in Johndrow et al., (2020) that several minibatch algorithms, including Maclaurin and Adams, (2014), but not those considered here, do not significantly improve performance for realistic problems like generalised linear models. Understanding this discrepancy may relate to the concept of “coresets,” as discussed in Huggins et al., (2016); Campbell and Broderick, (2018, 2019).
@@ A.3 Proof of Proposition 2
A.4 Algorithmic description of the idealised chain
Algorithm 7 idealised MH
1:  Initialise: Initial state $\theta_{0}$; number of iterations $T$; target distribution $\pi(\theta\mid x)$
@@ A.9 Proof of Proposition 5
It is useful to recall a few facts, which we provide proofs for completeness in supplementary A.11: for univariate Poisson random variables with parameter $\lambda_{1},\lambda_{2}$, their KL divergence equals $d_{\sf KL}(\mathsf{Poi}(\lambda_{1})||\mathsf{Poi}(\lambda_{2}))=\lambda_{1}\log(\lambda_{1}/\lambda_{2})+\lambda_{2}-\lambda_{1}$. Meanwhile, the KL divergence $d_{\sf KL}(\mathsf{Poi}(\lambda_{1})||\mathsf{Poi}(\lambda_{2}))$ satisfying $m\leq\lambda_{1}\leq\lambda_{2}\leq M$ is maximised when $\lambda_{1}=m$ and $\lambda_{2}=M$. Furthermore, KL divergence satisfies the ‘tensorization property’, i.e., $d_{\sf KL}(\otimes_{i}P_{i}||\otimes_{i}Q_{i})=\sum_{i}d_{\sf KL}(P_{i}||Q_{i})$, while $1-d_{\sf TV}(\otimes_{i}P_{i},\otimes_{i}Q_{i})\geq\prod_{i}(1-d_{\sf TV}(P_{i},Q_{i}))$.
@@
Let $\tilde{\lambda}_{i}(\theta):=\lambda M_{i}/L+\phi_{i}(\theta;x)$. We recognise that $\mathbb{P}_{\theta,\theta^{\prime}}=\otimes_{i=1}^{N}\mathsf{Poi}(\tilde{\lambda}_{i}(\theta))$ in PoissonMH. We have $1-d_{\sf TV}(\mathbb{P}_{\theta,\theta^{\prime}},\mathbb{P}_{\theta^{\prime},\theta})\geq\prod_{i}\left(1-d_{\sf TV}(\mathsf{Poi}(\tilde{\lambda}_{i}(\theta)),\mathsf{Poi}(\tilde{\lambda}_{i}(\theta^{\prime})))\right).$ For each term in the product, assuming without loss of generality that $\tilde{\lambda}_{i}(\theta^{\prime})>\tilde{\lambda}_{i}(\theta)$:
@@
We can also use the second part of Theorem 2, and calculate the symmetrised KL:
@@ A.10 Proof of Proposition 6
Let $\tilde{\phi}_{i}(\theta,\theta^{\prime}):=\lambda c_{i}/C+\phi_{i}(\theta,\theta^{\prime};x)$. We have $\mathbb{P}_{\theta,\theta^{\prime}}=\otimes_{i=1}^{N}\mathsf{Poi}(\tilde{\phi}_{i}(\theta,\theta^{\prime}))$ for TunaMH. We can calculate the symmetrised KL divergence
@@ A.11 Proof of auxiliary results
The proposition is equivalent to proving that the function $g(\lambda_{1},\lambda_{2}):=\lambda_{1}\log(\lambda_{1}/\lambda_{2})+\lambda_{2}-\lambda_{1}$, constrained within the region $m\leq\lambda_{1}\leq\lambda_{2}\leq M$, is maximised at $\lambda_{1}=m$ and $\lambda_{2}=M$. We first fix $\lambda_{1},$ and take derivative with respect to $\lambda_{2}$:
@@ A.12 Additional properties
Appendix B Implementation details of locally balanced PoissonMH
@@
Appendix C Additional comment on Algorithm 2 and 3
@@
Appendix D Further details of experiments in Section 5
D.1 Heterogeneous truncated Gaussian
@@
D.2 Robust Linear Regression
@@
D.3 Bayesian Logistic Regression on MNIST
@@
where the first inequality follows from Cauchy–Schwarz inequality and the second inequality comes from $|(h(x_{i}^{\top}\theta)-y_{i})|\leq 1$. Thus, we can take $M_{i}(\theta,\theta^{\prime})=\lVert\theta^{\prime}-\theta\rVert_{2}$ and $c_{i}=\lVert x_{i}\rVert_{2}$ for TunaMH and TunaMH–SGLD.
Appendix E Additional experiments
E.1 Additional experiments on 10-dimensional robust linear regression
Figure 6 compares the MSE over iteration count for all six algorithms in the robust linear regression example. It is clear that random-walk Metropolis outperforms PoissonMH, confirming our theoretical results in Section 4.1. However, PoissonMH has better performance in clock time due to its lower per-step cost. Poisson–MALA and MALA exhibit the best overall per-iteration performance, followed by HMC and Poisson–Barker. This indicates that our locally-balanced PoissonMH method can find proposals as effective as full batch methods while enjoying a very low per-step cost.
@@ E.1 Additional experiments on 10-dimensional robust linear regression
Figure 7: Clock-time MSE comparison for HMC with leapfrog steps $2,5,$ and $10$.
E.2 50-dimensional robust linear regression
We compare Poisson-Barker, Poisson-MALA, PoissonMH, random-walk Metropolis, MALA, and HMC in a 50-dimensional robust linear regression example. All setups are the same as in Section 5.2, except for $d=50$, and the number of leapfrog steps in HMC is 5.
@@
ESS/s: (Min, Median, Max) Method acceptance rate=0.25 acceptance rate=0.4 acceptance rate=0.55 Best MH (0.09, 0.11, 0.14) (0.08, 0.10, 0.12) (0.06, 0.07, 0.08) (0.09, 0.11, 0.14) MALA (0.40, 0.49, 0.55) (0.77, 0.86, 0.97) (0.82, 0.89, 0.97) (0.82, 0.89, 0.97) HMC-5 (0.04, 0.04, 0.05) (0.04, 0.04, 0.05) (0.03, 0.03, 0.04) (0.04, 0.04, 0.05) PoissonMH (2.12, 2.18, 2.30) (1.85, 1.95, 2.03) (1.36, 1.43, 1.48) (2.12, 2.18, 2.30) PoissonMH-Barker (4.99, 5.25, 5.62) (6.55, 7.05, 7.57) (6.99, 7.37, 7.57) (6.99, 7.37, 7.57) PoissonMH-MALA (8.35, 8.82, 9.16) (13.34, 14.36, 14.81) (15.89, 16.15, 16.58) (15.89, 16.15, 16.58)
@@
Figure 10: Clock-time MSE comparison for HMC with leapfrog steps $2,5,$ and $10$.
E.3 Additional experiments on Bayesian Logistic Regression: 3s and 5s
@@
Figure 12 shows a comparison of all six algorithms when each is optimally tuned. From the left panel, it is clear that Tuna–SGLD converges fastest among all algorithms, followed by TunaMH and random-walk Metropolis. The right panel indicates that MALA exhibits superior iteration-wise convergence compared to Tuna–SGLD, while random-walk Metropolis demonstrates better iteration-wise convergence than TunaMH. This observation supports our theoretical findings in Section 4.1.
@@
Figure 12: Test accuracy comparison when each algorithm is optimally tuned
E.4 Bayesian Logistic Regression: 7s and 9s
Similar to Section 5.3, we test TunaMH, TunaMH–SGLD, random–walk Metropolis, MALA and HMC on Bayesian logistic regression task to classify 7s and 9s in the MNIST data set. The training set contains 12,214 samples, and the test set contains 2,037 samples. We adopt all implementation details from Section 5.3 and report the results in Figure 13. We can draw similar conclusions to Section 5.3 that our proposed minibatch gradient-based algorithm performs better than its original minibatch non-gradient version in all step size configurations. All minibatch methods still significantly improve the performances of full batch methods.

### v3

@@ 6 Future directions
Another intriguing question is the source of performance gains in minibatch MCMC algorithms. Empirically, these algorithms often achieve ESS/s values 10-100 times greater than MCMC algorithms running on the entire dataset. However, it is proven in Johndrow et al., (2020) that several minibatch algorithms, including Maclaurin and Adams, (2014), but not those considered here, do not significantly improve performance for realistic problems like generalized linear models. Understanding this discrepancy may relate to the concept of “coresets,” as discussed in Huggins et al., (2016); Campbell and Broderick, (2018, 2019).
@@ A.3 Proof of Proposition 2
A.4 Algorithmic description of the idealized chain
Algorithm 7 idealized MH
1:  Initialize: Initial state $\theta_{0}$; number of iterations $T$; target distribution $\pi(\theta\mid x)$
@@ A.9 Proof of Proposition 5
It is useful to recall a few facts, which we provide proofs for completeness in supplementary A.11: for univariate Poisson random variables with parameter $\lambda_{1},\lambda_{2}$, their KL divergence equals $d_{\sf KL}(\mathsf{Poi}(\lambda_{1})||\mathsf{Poi}(\lambda_{2}))=\lambda_{1}\log(\lambda_{1}/\lambda_{2})+\lambda_{2}-\lambda_{1}$. Meanwhile, the KL divergence $d_{\sf KL}(\mathsf{Poi}(\lambda_{1})||\mathsf{Poi}(\lambda_{2}))$ satisfying $m\leq\lambda_{1}\leq\lambda_{2}\leq M$ is maximized when $\lambda_{1}=m$ and $\lambda_{2}=M$. Furthermore, KL divergence satisfies the ‘tensorization property’, i.e., $d_{\sf KL}(\otimes_{i}P_{i}||\otimes_{i}Q_{i})=\sum_{i}d_{\sf KL}(P_{i}||Q_{i})$, while $1-d_{\sf TV}(\otimes_{i}P_{i},\otimes_{i}Q_{i})\geq\prod_{i}(1-d_{\sf TV}(P_{i},Q_{i}))$.
@@
Let $\tilde{\lambda}_{i}(\theta):=\lambda M_{i}/L+\phi_{i}(\theta;x)$. We recognize that $\mathbb{P}_{\theta,\theta^{\prime}}=\otimes_{i=1}^{N}\mathsf{Poi}(\tilde{\lambda}_{i}(\theta))$ in PoissonMH. We have $1-d_{\sf TV}(\mathbb{P}_{\theta,\theta^{\prime}},\mathbb{P}_{\theta^{\prime},\theta})\geq\prod_{i}\left(1-d_{\sf TV}(\mathsf{Poi}(\tilde{\lambda}_{i}(\theta)),\mathsf{Poi}(\tilde{\lambda}_{i}(\theta^{\prime})))\right).$ For each term in the product, assuming without loss of generality that $\tilde{\lambda}_{i}(\theta^{\prime})>\tilde{\lambda}_{i}(\theta)$:
@@
We can also use the second part of Theorem 2, and calculate the symmetrized KL:
@@ A.10 Proof of Proposition 6
Let $\tilde{\phi}_{i}(\theta,\theta^{\prime}):=\lambda c_{i}/C+\phi_{i}(\theta,\theta^{\prime};x)$. We have $\mathbb{P}_{\theta,\theta^{\prime}}=\otimes_{i=1}^{N}\mathsf{Poi}(\tilde{\phi}_{i}(\theta,\theta^{\prime}))$ for TunaMH. We can calculate the symmetrized KL divergence
@@ A.11 Proof of auxiliary results
The proposition is equivalent to proving that the function $g(\lambda_{1},\lambda_{2}):=\lambda_{1}\log(\lambda_{1}/\lambda_{2})+\lambda_{2}-\lambda_{1}$, constrained within the region $m\leq\lambda_{1}\leq\lambda_{2}\leq M$, is maximized at $\lambda_{1}=m$ and $\lambda_{2}=M$. We first fix $\lambda_{1},$ and take derivative with respect to $\lambda_{2}$:
@@ A.12 Additional properties
Appendix B Implementation details of Poisson Minibatching
We provide the implementation details for sampling the minibatch data in Algorithm 8 (for Algorithm 2, 5), and Algorithm 9 (for Algorithm 3, 6). The goal is to demonstrate that this minibatch sampling step can be implemented at a much lower cost than its naïve implementation (and also lower than full batch MCMC). All the material here is not new, as this subsampling trick has been extensively discussed in Zhang and De Sa, (2019); Zhang et al., (2020). Since the same subsampling method is used in our new algorithm (Algorithm 5, 6), we include their details here for concreteness.
Algorithm 8 Minibatch Sampling in PoissonMH and Locally Balanced PoissonMH
  $\triangleright$ One time pre-computation
  • set $\Lambda_{i}:=\frac{\lambda M_{i}}{L}+M_{i}$ and $\Lambda:=\sum_{i=1}^{N}\Lambda_{i}$,
  • construct an alias table for the categorical distribution on $\{1,2,\cdots,N\}$: $\texttt{Categorical}(p_{1},p_{2},\cdots,p_{N})$, where $p_{i}=\Lambda_{i}/\Lambda$
  $\triangleright$ Poisson minibatch sampling (Step 5 of Algorithm 2, Step 3 of Algorithm 5)
  Current position $\theta$
  Initialize $s_{i}=0$ for $i\in\{1,2,\ldots,N\}$
  sample $B\sim\mathsf{Poi}(\Lambda)$
  for $b=1$ to $B$ do
     sample $i_{b}\sim\texttt{Categorical}(p_{1},p_{2},\cdots,p_{N})$
     compute $\phi_{i_{b}}(\theta;x)$
     with probability $\frac{\frac{\lambda M_{i_{b}}}{L}+\phi_{i_{b}}(\theta;x)}{\frac{\lambda M_{i_{b}}}{L}+M_{i_{b}}}$, set $s_{i_{b}}\leftarrow s_{i_{b}}+1$
  end for
Algorithm 9 Minibatch Sampling in TunaMH and Tuna–SGLD
  $\triangleright$ One time pre-computation before MCMC iterations
  • construct an alias table for the categorical distribution on $\{1,2,\cdots,N\}$: $\texttt{Categorical}(p_{1},p_{2},\cdots,p_{N})$, where $p_{i}=c_{i}/C$
  $\triangleright$ Poisson minibatch sampling (Step 6 of Algorithm 3, Step 3 of Algorithm 6)
  Current position $\theta$ and proposed state $\theta^{\prime}$
  Initialize $s_{i}=0$ for $i\in\{1,2,\ldots,N\}$
  sample $B\sim\mathsf{Poi}(\lambda+CM(\theta,\theta^{\prime}))$
  for $b=1$ to $B$ do
     sample $i_{b}\sim\texttt{Categorical}(p_{1},p_{2},\cdots,p_{N})$
     compute $\phi_{i_{b}}(\theta,\theta^{\prime};x)$
     with probability $\frac{\lambda c_{i_{b}}+C\phi_{i_{b}}(\theta,\theta^{\prime};x)}{\lambda c_{i_{b}}+Cc_{i_{b}}M(\theta,\theta^{\prime})}$, set $s_{i_{b}}\leftarrow s_{i_{b}}+1$
  end for
Both algorithms require pre-computation that only needs to be performed once prior to the MCMC iterations. This computation constructs an ‘alias table’ for sampling the categorical distribution on $\{1,2,\cdots,N\}$. With this table, drawing a sample from this categorical distribution requires only $\mathcal{O}(1)$ time. See part A and part B in Section III of Vose, (1991) for details.
The computational cost of generating Poisson variables at each MCMC step is then $\mathcal{O}(B)$. The expected computational cost will be:
$\mathbb{E}[B]=\begin{cases}\lambda+L&\text{for PoissonMH}\\ \lambda+CM(\theta,\theta^{\prime})&\text{for TunaMH}\end{cases}$
Here $\lambda$ is our tuning parameter, and $L$ and $C$ are much smaller than $N$ in the examples considered here. Therefore the per-iteration cost is of these minibatch MCMC algorithm is much smaller than standard MCMC.
Next, we verify that Algorithm 8 indeed generate independent Poisson variables required by Step 5 of PoissonMH (Algorithm 2). The validity proof of Algorithm 9 is similar. For PoissonMH, we can view the generating process of $\{s_{i}\}_{i=1}^{N}$ as the following urn model: 1. Generates $\mathsf{Poi}(\Lambda)$ balls; 2. Throw each ball to $N$ urns independently, with probability $p_{1},p_{2},\ldots,p_{N}$ for the first, second, …, last urn. 3. For every ball in urn $i$ (if any), keep it with probability $\frac{\frac{\lambda M_{i_{b}}}{L}+\phi_{i_{b}}(\theta_{t};x)}{\frac{\lambda M_{i_{b}}}{L}+M_{i_{b}}}$ and discard it otherwise. We can verify $s_{1},s_{2},\ldots,s_{N}$ are the number of balls in urn $1,2,\ldots,N$ respectively.
On the other hand, step 1 - 3 are precisely the Poisson thinning process. Standard Poisson calculation shows $s_{1},s_{2},\ldots,s_{N}$ are independent Poisson distribution, and the Poisson parameter of $s_{i}$ equals $B\times p_{i}\times\frac{\frac{\lambda M_{i}}{L}+\phi_{i}(\theta_{t};x)}{\frac{\lambda M_{i}}{L}+M_{i}}$. In our case, $p_{i}=\Lambda_{i}/\Lambda$, and $\Lambda_{i}=\lambda M_{i}/L+M_{i}$. Therefore $s_{i}\sim\mathsf{Poi}(\lambda M_{i}/L+\phi_{i}(\theta,x))$.
As a result, Algorithm 8 gives the required Poisson variables by PoissonMH. For TunaMH, the proof is similar and is omitted here.
Appendix C Implementation details of locally balanced PoissonMH
@@
Appendix D Additional comment on Algorithm 2 and 3
@@
Appendix E Further details of experiments in Section 5
E.1 Heterogeneous truncated Gaussian
@@
E.2 Robust Linear Regression
@@
E.3 Bayesian Logistic Regression on MNIST
@@
where the first inequality follows from Cauchy–Schwarz inequality and the second inequality comes from $|(h(x_{i}^{\top}\theta)-y_{i})|\leq 1$. Thus, we can take $M_{i}(\theta,\theta^{\prime})=\lVert\theta^{\prime}-\theta\rVert_{2}$ and $c_{i}=\lVert x_{i}\rVert_{2}$ for TunaMH and Tuna–SGLD.
Appendix F Additional experiments
F.1 Additional experiments on 10-dimensional robust linear regression
MALA and Poisson–MALA exhibit the best overall per-iteration performance, followed by HMC, Barker and Poisson–Barker. This indicates that our locally-balanced PoissonMH method can find proposals as effective as full batch methods while enjoying a very low per-step cost.
@@ F.1 Additional experiments on 10-dimensional robust linear regression
Figure 7: Clock-time MSE comparison for HMC with leapfrog steps $2,5,$ and $10$. All results are averaged over 10 runs.
F.2 50-dimensional robust linear regression
We compare Poisson-Barker, Poisson-MALA, PoissonMH, random-walk Metropolis, MALA, Barker and HMC in a 50-dimensional robust linear regression example. All setups are the same as in Section 5.2, except for $d=50$, and the number of leapfrog steps in HMC is 5.
@@
ESS/s: (Min, Median, Max) Method acceptance rate=0.25 acceptance rate=0.4 acceptance rate=0.55 Best MH (0.08, 0.10, 0.13) (0.07, 0.08, 0.12) (0.06, 0.07, 0.08) (0.00, 0.10, 0.12) MALA (0.39, 0.48, 0.56) (0.63, 0.72, 0.81) (0.73, 0.83, 0.91) (0.73, 0.83, 0.91) Barker (0.17, 0.21, 0.25) (0.23, 0.26, 0.29) (0.23, 0.28, 0.31) (0.23, 0.28, 0.31) HMC-5 (0.03, 0.04, 0.05) (0.03, 0.03, 0.04) (0.02, 0.03, 0.04) (0.03, 0.04, 0.05) PoissonMH (1.91, 2.04, 2.21) (1.57, 1.66, 1.83) (1.24, 1.34, 1.44) (1.91, 2.04, 2.21) PoissonMH-Barker (4.18, 4.54, 4.78) (5.47, 5.85, 6.29) (6.07, 6.51, 6.84) (6.07, 6.51, 6.84) PoissonMH-MALA (6.76, 7.45, 8.11) (11.00, 11.81, 12.40) (13.92, 14.91, 15.58) (13.92, 14.91, 15.58)
@@
Figure 10: Clock-time MSE comparison for HMC with leapfrog steps $2,5,$ and $10$. All results are averaged over 10 runs.
F.3 Additional experiments on Bayesian Logistic Regression: 3s and 5s
@@
Figure 12 shows a comparison of all seven algorithms when each is optimally tuned. From the left panel, it is clear that Tuna–SGLD converges fastest among all algorithms, followed by TunaMH and random-walk Metropolis. The right panel indicates that MALA exhibits superior iteration-wise convergence compared to Tuna–SGLD, while random-walk Metropolis demonstrates better iteration-wise convergence than TunaMH. This observation supports our theoretical findings in Section 4.1.
@@
Figure 12: Test accuracy comparison when each algorithm is optimally tuned.
F.4 Bayesian Logistic Regression: 7s and 9s
Similar to Section 5.3, we test TunaMH, Tuna–SGLD, random–walk Metropolis, MALA, Barker and HMC on Bayesian logistic regression task to classify 7s and 9s in the MNIST data set. The training set contains 12,214 samples, and the test set contains 2,037 samples. We adopt all implementation details from Section 5.3 and report the results in Figure 13. We can draw similar conclusions to Section 5.3 that our proposed minibatch gradient-based algorithm performs better than its original minibatch non-gradient version in all step size configurations. All minibatch methods still significantly improve the performances of full batch methods.